A wire 75 cm long carrying a current of 6.0 A is at right angles to a uniform magnetic field. The magnitude of the force acting on the wire is 0.60 N. What is the strength of the magnetic field ?
1
Expert's answer
2011-08-25T13:12:59-0400
According to the Ampere's law:


F=B*I*L*sinA.,
where
F - the Ampere's force,
B - strength of the magnetic field,
I - amperage in a wire,
L - length of a wire,
A - angle between the amperage and the vector of magnetic field


As the wire is at the right angles, we will get that:


sinA=1,
B=F/(I*L)
B=0.6/(6*0.75)=0.13 (tesla)


The strength of the magnetic field is 0.13 tesla.